Sulfur dioxide (SO2
), a gas that plays a central role in the formation of acid rain, is found in the exhaust of
automobiles and power plants. Consider a 1.53-L sample of gaseous SO2 at a pressure of 5.6x103 Pa. If the
pressure is changed to 1.5x104 Pa at a constant temperature, what will be the new volume of the gas?
1
Expert's answer
2020-10-15T03:22:07-0400

the temperature does not change. hence we do so on the basis of the Boyle Marriott law. 

Boyle Marriott's law: the pressure of a gas at constant temperature is inversely proportional to its volume. that is,

P1 / P2 = V 2/ V1

where, P1 —is the initial pressure, P2 —is the next pressure, V1 —is the initial volume, and V2—is the next volume

according to the circumstances of the case

V1=1.53 L

P1=5.6*103 Pa

P2=1.5*104 Pa = 15*103 Pa

V2=?

V2=P1*V1/P2= 5.6*103* 1.53/15*103= 0.5712 L or 571.2 ml

ANSWER: 0.5712 L or 571.2ml